# Autonomation, not Automation: Activities and Needs of Fact-checkers as a Basis for Designing Human-Centered AI Systems

### v2

@@ 4.3. Communication and avoiding duplication
Communication within the teams in fact-checking organizations (Fig. 6) is not usually considered an issue, whereas fact-checkers utilize traditional communication channels together with freely available tools (e.g. Google Docs, instant messaging services). Nevertheless, inter-organizational and preferably international communication is very important for fact-checkers as they need 1) to be aware of the fact-checks that are being verified or published, and 2) for the coordination with other fact-checkers worldwide.
@@
And it is the coordination of fact-checkers with other organizations globally (Fig. 6) that our respondents considered problematic. Even though fact-checkers search for fact-checks from other organizations, they have been aware of just about some of the published fact-checks. Communication or some autonomated identification of misinformation among the fact-checking and journalistic organizations globally is also important to identify local misinformation in one’s country quicker, as this fact-checker (P2) responded:
Collaboration between fact-checking organizations is recognized as essential throughout the world to share tips and technical knowledge (Micallef et al., 2022). However, Central-European fact-checkers working in small NGOs do not share such tips. In some regions such as, for example, the Balkan region or Latin America, it is common practice to actively share such tips in mailing lists, regional forums, or WhatsApp groups related to fact-checking (Micallef et al., 2022; Juneja and Mitra, 2022). Some, like Factify even set up e-learning platforms with video tutorials just for fact-checkers  (Juneja and Mitra, 2022). Collaboration, building coalitions with other fact-checkers and shared training are the fields that need to be improved in the region to strengthen the positions of fact-checkers.
@@ 4.4. Verification of content credibility and veracity
The evaluation of the claims is considered an activity that requires human judgment. In our region as well as throughout the world (noa, 2020; Micallef et al., 2022; Juneja and Mitra, 2022), this activity requires communication with people. Fact-checkers regard domain experts as important sources of information. Sometimes, fact-checkers confront authors of the checked content, asking them to back up their claims or respond to fact-checks. Communication also involves requesting or clarifying data from government officials and press agencies, which slows down the process significantly as these officials are very careful about their statements. The complexity of the task can be considered as an obstacle to automation.
Nevertheless, another time and labor-intensive stage, viewed by fact-checkers as a potential candidate for autonomation, is evidence retrieval. Basically, it involves the acquisition of primary sources to achieve the maximum possible objectivity (Fig. 7). Our focus on human-information interaction enabled us to go deeper into the specific resources that the fact-checkers use. The textual sources involve mostly official data (as statistics), official information on the websites of governments, private companies, and NGOs, or original research studies. A fact-checker (P3) describes the process:
“We have to find the right source for the data, such as in the Polish statistical bureau, central controller’s office, Eurostat… If the claim is not about data, we try to find facts or proof since as journalists, we can be sued for everything we write… It is not just debunking whether [the claim] is true or not… Lots of claims are manipulative or partially true, … Very often we use experts (such as in finance, energy, or climate) to evaluate  information in statements - including implicit ones… The media is not our sources, they are good for background or if someone said something in the media.”
@@
InVid tool in the WeVerify plugin was seen as very relevant for our respondents as it decomposes the videos into keyframes (images) that are possible to be reverse-searched. Nevertheless, as some misinformation is very complex, fact-checkers need to do more than reverse-search the images. Interestingly, participants in the study of (Micallef et al., 2022), who reside mostly outside of Europe, do not report being aware of these tools.
Claim verification is seen by our respondents as an “intuitive process” that results from the long-term practice of fact-checkers, often journalists. Fact-checkers prevailingly stated that they notice questionable sources “at first sight”. However, one respondent shared his credibility indicators for the evaluation of sources, namely: non-existent author, problematic source (in the past), factuality, distortions, biases, objectivity, out-of-the-context use, sentiment and argumentation fallacies.
A “contextualization stage” was mentioned (Micallef et al., 2022), where most fact-checkers of their study looked into the evolution of the claim from its origin to the present state to help readers understand their conclusions. Nonetheless, this process is very challenging, especially in foreign languages. One of our interviewees mentioned the problematic tracing of such a history, particularly on social media, where the search capabilities are very limited.
The verification process is considered very transparent, as the websites of fact-checking organizations often explain how the judgment was reached, attaching the sources as well (often their archived versions). Besides that, it is very rigorous. As reported, none of the fact-checks is published without consulting it with at least one editor. Participants from Africa and Oceania have even longer peer review processes than Western countries (Micallef et al., 2022).
@@ 4.5. Dissemination of fact-checks
Because of capacity reasons, dissemination of fact-checks mostly takes place where the misinformation spreads - on social media, mostly Facebook (Fig. 8). Fact-checks are published on the websites of the organizations. In connection to web publishing, Central European fact-checkers frequently report limited technological capabilities in their organizations (unlike some fact-checkers throughout the world (noa, 2020; Micallef et al., 2022)). These prevent them from full exploitation of search engine optimization (SEO), ClaimReview, appearing in Google News, or paid ads on Google as noted by P5:
@@
Our respondents from small NGOs also struggle with personal (marketing and design) capacities to be able to make their fact-checks more appealing and popular by including visuals (e.g. infographics, videos, comic stripes). This is a common practice for copy editors in some fact-checking organizations  (Juneja and Mitra, 2022). Nevertheless, fact-checkers communicate with the media, as P6 claimed:
@@ 5.1. The problems of European fact-checkers
In general, we can conclude that the validation survey supports and extends the findings of the in-depth interviews. Regarding the differences, the survey showed a higher urgency to autonomize alerts and user tips from instant messaging services and filtering the monitored outlets. The coordination with other fact-checkers and the analysis of the impact of the fact-checks was perceived as an issue with a lower urgency than it seemed from the interviews. Thus, these problems are considered specific to the region and need to be addressed locally.
@@
According to the results form interviews, the verification of the truthfullness of the claim is not perceived as an issue. However, some parts of the verification process take most of the valuable time of fact-checkers and would urgently need some AI support. The validation survey confirmed that one of the biggest problems relates to searching for the sources of evidence for verification, particularly in hard-to-find official documents, videos, and statistics. This is especially true for cases where data needs to be integrated, as one fact-checker (P10) noted:
@@ 5.2.1. Differences between the problems of Eastern and Western European fact-checkers
The data do not show many differences between the problems of Eastern and Western European fact-checkers (Fig. 10). However, searching within the existing fact-checks (both their own and from other fact-checking organizations and in potentially different languages) is considered a more urgent problem to solve by Eastern European fact-checkers. This identified need is relevant to address, as the existing fact-checks are an important source of evidence for these professionals. The need for a tool that would help to autonomate this part of the process was confirmed by a fact-checker (P9) in a comment of the survey:
“Would totally be a great help. Compiling all the IFCN signatories’ outputs into one big (keyword-driven) database? A fact-checker’s dream. Although, it is fair to say that Google Search supplements a lot of this proposition - although it’s always difficult with other languages than English - relevant now for Ukrainian.“
@@
Russian trolls know that in order to be successful, they have to go through the west (Germany, for example), because Poles do not like Russian sources… Also, sometimes what is popular abroad - mostly in Czech or English, such as vaccines - is popular in Poland as well.”
In contrast, western European fact checkers focus more on misinformation modalities beyond text and platforms beyond Facebook. This is demonstrated by their pressing need to autonomate alerts from instant messaging services, such as WhatsApp, possibly Telegram, and a more urgent need to search for image manipulations. The less perceived urgency of these needs by Eastern European respondents can be explained by the lack of capacity that allows them to focus mainly on text and Facebook, as mentioned during some of the interviews.
@@ 6. Implications for AI research and AI-based tools
To answer RQ3, we identified implications and opportunities for research and development of AI tools that would support fact-checkers in fulfilling their tasks. For clarity, we mapped the stages of fact-checking process, the most urgent problems of fact-checkers (resulting from our validation survey) with the corresponding implications for AI support (Fig. 13).
Figure 13. Mapping of the top 10 most important problems of fact-checkers with the corresponding stages of the fact-checking process and implications for AI support
@@
Our proposed implications for AI support come from the current state-of-the-art research and technical possibilities in the area of AI – machine learning (ML) and NLP (including the latest development of generative large language models), and also from available datasets and tools and are consistent with the work by (Nakov et al., 2021a). Compared to (Nakov et al., 2021a), we add three more AI tasks: (1) check-worthy document detection, (2) mapping of the existing fact-checks to additional (newly appearing) online content, and (3) fact-check summarization and personalization.
At the same time, we recognize that some of the most important problems do not require AI-based solutions. In such cases, the implementation of suitable tools represents mainly an engineering challenge (e.g., how to implement a tool that fact-checkers can use to insert ClaimReview schema into their fact-checks), which is out of the scope of this work.
@@ 6.1. Implications for Check-worthy document detection
The first AI task aims to support filtering the check-worthy documents (that is, news articles, blog posts, or even posts on social media) from monitored sources. It can be defined as follows: given an input set of documents, detect such documents that are worth fact-checking, and thus they should receive fact-checkers’ attention (e.g., they contain factual claims, they are potentially impactful and harmful to society, etc.). This task can be addressed either as a classification (a document is or is not check-worthy) or a ranking problem (prioritizing the most check-worthy documents). It should be addressed with information available at document creation time (i.e., without relying on user feedback that appears later). Given that false information spreads faster than true information (Vosoughi et al., 2018), the impact of a fact-check can be limited once the post has already become viral. On the other hand, early fact-checks can limit the spread of false information and can thus serve as a way of pre-bunking, which has already been shown to be effective in existing studies (Roozenbeek et al., 2022).
@@
We argue that this task should generally precede the task of check-worthy claims detection (extraction) to limit and prioritize the amount of potential misinformation content the fact-checkers need to examine. Current systems (e.g., a fact-checking tool used within Meta’s Third-Party Fact-Checking Program) often prioritize the virality of a post as noted by the fact-checkers in our as well as in the previous studies (noa, 2020; Dierickx, 2022; Micallef et al., 2022). In research work, check-worthiness detection is originally related to claims (i.e., identification of particular sentences, typically in political debates). Nevertheless, starting in 2020, the CheckThat! Lab introduced the detection of check-worthy tweets (Hasanain and Elsayed, 2022). Similarly, (Alam et al., 2021) introduced a holistic approach to annotate 7 features reflecting the check-worthiness of COVID-19 related tweets. Although tweets are natively short texts containing just a few sentences, we would like to emphasize distinction with check-worthy claim detection (see below) and we consider this as the first step towards check-worthy document detection. At the moment, we are not aware of any works on check-worthiness detection for longer pieces of text (e.g., for social media posts from other platforms that do not pose such strict length limitations, or even for whole news articles/blogs) or for multimodal content. Such extension of current research works and construction of a suitable dataset represents the possible next step.
@@
Check-worthy document detection may be considered as related to a more generic task of credibility indicators assessment. Credibility indicators (e.g., credibility signals proposed by W3C Credible Web Community Group) would similarly help fact-checkers (as well as other media professionals) to pre-screen the document and decide whether to proceed with the manual in-depth investigation to determine the necessity to fact-check it. Such a tool would speed up fact-checkers’ comprehension of the online content. To detect such indicators, a wide variety of techniques may be used, from a simple lookup in whitelists/blacklists, through the automatic check of pre-defined credibility criteria (e.g., presence of an author’s name, the known editorial board of a newspaper, etc.) up to advanced ML/NLP models. These would, for example, classify typographical and stylistic characteristics reflecting psychological features influencing the reader’s sentiment, detect the use of logical fallacies, or classify the leaning (bias) of the text (Martino et al., 2020; Dhoju et al., 2019). The detected check-worthiness of documents can serve as one of the indicators if the detection also considers their potential harmfulness or their expected impact on society (see, e.g., (Alam et al., 2021)), which are the aspects usually considered by the fact-checkers (see Fig. 5).
@@ 6.2. Implications for Check-worthy claims detection
While the task of check-worthy document detection (and the related task of credibility signals extraction) has attracted researchers’ attention only recently, the check-worthy claims detection has been addressed by the researchers for a longer period and several established approaches, as well as datasets, already exist (see, e.g., (Hassan et al., 2017)). This task can be defined as follows: given an input sentence (or a set of multiple sentences, for example, from a political debate), detect such sentences that are worthy of fact-checking [^1] While the definition of a check-worthy claim differs across literature, it is usually a factual statement, that the general public would be interested in knowing, whether it is true or not.. Most of the existing approaches have so far focused on text using transformer-based language models, such as BERT (Devlin et al., 2019). However, there is a growing need for multimodal approaches, since claims can be accompanied by images, they can themselves appear in images (e.g., a screenshot of a social media post) or images can modify/change their meanings. There are already the first multimodal datasets for claim detection (see, e.g., (Zlatkova et al., 2019; Cheema et al., 2022)).

### v3

@@ 4.3. Communication and avoiding duplication
Communication within fact-checking teams (Fig. 6) is generally not perceived as problematic. Teams typically rely on traditional communication channels and freely available tools such as Google Docs and instant messaging services. However, interorganizational—and especially international—communication is considered crucial, as fact-checkers need: 1) to stay informed about fact-checks currently being verified or published, and 2) to coordinate efforts with colleagues worldwide.
It is precisely this global coordination (Fig. 6) that respondents identified as a challenge. While fact-checkers actively search for fact-checks from other organizations, they are usually aware of only a fraction of those published. Improving communication or implementing autonomated systems for identifying misinformation across fact-checking and journalistic organizations could accelerate also the identification of local misinformation, as one fact-checker (P2) explained:
@@
Collaboration among fact-checking organizations is widely recognized as essential for sharing tips and technical expertise (Micallef et al., 2022). In some regions, such as the Balkans or Latin America, it is common practice to exchange information through mailing lists, regional forums, or WhatsApp groups dedicated to fact-checking (Micallef et al., 2022; Juneja and Mitra, 2022). Some initiatives, such as Factify, have even developed e-learning platforms with video tutorials exclusively for fact-checkers (Juneja and Mitra, 2022). By contrast, Central European fact-checkers—often working in small NGOs—rarely engage in such exchanges. Strengthening collaboration, building coalitions, and developing shared training opportunities are key areas for improvement in the region to enhance the capacity and influence of fact-checkers.
@@ 4.4. Verification of content credibility and veracity
Content evaluation is widely regarded as an activity that requires human judgment. Both in our region and globally (noa, 2020; Micallef et al., 2022; Juneja and Mitra, 2022), this process depends heavily on communication with people. Fact-checkers view domain experts as valuable sources of information, and at times, they contact the authors of disputed content to request evidence or responses to fact-checks. They also engage with government officials and press agencies to request or clarify data—a step that can significantly delay the process, as such officials tend to be highly cautious in their statements. This reliance on interpersonal communication and expert input underscores the complexity of content evaluation and represents a key obstacle to its automation.
Nevertheless, while evaluation itself resists full automation, fact-checkers view evidence retrieval—a time- and labor-intensive stage—as a promising candidate for technological support and partial automation. Evidence retrieval involves the acquisition of primary sources to achieve the maximum possible objectivity (Fig. 7). Our focus on human-information interaction enabled us to go deeper into the specific resources that the fact-checkers use for evidence retrieval. The textual sources involve mostly official data (as statistics), official information on the websites of governments, private companies, and NGOs, or original research studies. A fact-checker (P3) describes the process:
“We have to find the right source for the data, such as in the Polish statistical bureau, central controller’s office, Eurostat… If the claim is not about data, we try to find facts or proof since as journalists, we can be sued for everything we write… It is not just debunking whether [the claim] is true or not… Lots of claims are manipulative or partially true, … Very often we use experts (such as in finance, energy, or climate) to evaluate  information in statements – including implicit ones… The media is not our sources, they are good for background or if someone said something in the media.”
@@
InVid tool in the Verification plugin was seen as very relevant for our respondents as it decomposes the videos into keyframes (images) that are possible to be reverse-searched. Nevertheless, as some misinformation is very complex, fact-checkers need to do more than reverse-search the images. Interestingly, participants in the study of (Micallef et al., 2022), who reside mostly outside of Europe, do not report being aware of these tools.
Content verification is seen by our respondents as an “intuitive process” that results from the long-term practice of fact-checkers, often journalists. Fact-checkers prevailingly stated that they notice questionable sources “at first sight”. However, one respondent shared his credibility indicators for the evaluation of sources, namely: non-existent author, problematic source (in the past), factuality, distortions, biases, objectivity, out-of-the-context use, sentiment and argumentation fallacies.
In previous research (Micallef et al., 2022) a “contextualization stage” was mentioned, where most fact-checkers looked into the evolution of the claim from its origin to the present state to help readers understand their conclusions. This process is very challenging, especially in foreign languages. One of our interviewees mentioned the problematic tracing of such a history, particularly on social media, where the search capabilities are very limited.
Fact-checkers’ verification processes are generally very transparent, as the websites of fact-checking organizations often explain how the judgment was reached, attaching the sources as well (often their archived versions). Besides that, it is very rigorous. As reported, none of the fact-checks is published without consulting it with at least one editor. Participants from Africa and Oceania have even longer peer review processes than Western countries (Micallef et al., 2022).
@@ 4.5. Dissemination of fact-checks
Because of capacity reasons, dissemination of fact-checks mostly takes place where the misinformation spreads – on social media, mostly Facebook (Fig. 8). Fact-checks are published on the websites of the organizations. In connection to web publishing, Central European fact-checkers frequently report limited technological capabilities in their organizations (unlike some fact-checkers throughout the world (noa, 2020; Micallef et al., 2022)). These prevent them from full exploitation of search engine optimization (SEO), ClaimReview, appearing in Google News, or paid ads on Google as noted by P5:
@@
Our respondents from small NGOs also struggle with personal (marketing and design) capacities to be able to make their fact-checks more appealing and popular by including visuals (e.g., infographics, videos, comic stripes). This is a common practice for copy editors in some fact-checking organizations  (Juneja and Mitra, 2022). Nevertheless, fact-checkers communicate with the media, as P6 claimed:
@@ 5.1. The problems of European fact-checkers
In general, we can conclude that the validation survey supports and extends the findings of the in-depth interviews. Regarding the differences, the survey showed a higher urgency to autonomate alerts and user tips from instant messaging services and filtering the monitored outlets. The coordination with other fact-checkers and the analysis of the impact of the fact-checks was perceived as an issue with a lower urgency than it seemed from the interviews. Thus, these problems are considered specific to the region and need to be addressed locally.
@@
According to the results from interviews, the verification of the truthfullness of the claim is not perceived as an issue. However, some parts of the verification process take most of the valuable time of fact-checkers and would urgently need some AI support. The validation survey confirmed that one of the biggest problems relates to searching for the sources of evidence for verification, particularly in hard-to-find official documents, videos, and statistics. This is especially true for cases where data needs to be integrated, as one fact-checker (P10) noted:
@@ 5.2.1. Differences between the problems of Eastern and Western European fact-checkers
The data do not show many differences between the problems of Eastern and Western European fact-checkers (Fig. 10). However, searching within the existing fact-checks (both their own and from other fact-checking organizations and in potentially different languages) is considered a more urgent problem to solve by Eastern European fact-checkers. This recognized need is relevant to address, as the existing fact-checks are an important source of evidence for these professionals. The need for a tool that would help to autonomate this part of the process was confirmed by a fact-checker (P9) in a comment of the survey:
“Would totally be a great help. Compiling all the IFCN signatories’ outputs into one big (keyword-driven) database? A fact-checker’s dream. Although, it is fair to say that Google Search supplements a lot of this proposition – although it’s always difficult with other languages than English – relevant now for Ukrainian.“
@@
Russian trolls know that in order to be successful, they have to go through the West (Germany, for example), because Poles do not like Russian sources… Also, sometimes what is popular abroad – mostly in Czech or English, such as vaccines – is popular in Poland as well.”
In contrast, Western European fact-checkers focus more on misinformation modalities beyond text and platforms beyond Facebook. This is demonstrated by their pressing need to autonomate alerts from instant messaging services, such as WhatsApp, possibly Telegram, and a more urgent need to search for image manipulations. The less perceived urgency of these needs by Eastern European respondents can be explained by the lack of capacity that allows them to focus mainly on text and Facebook, as mentioned during some of the interviews.
@@ 6. Implications for AI research and AI-based tools
To answer RQ3, we identified implications and opportunities for research and development of AI tools that would support fact-checkers in fulfilling their tasks. For clarity, we mapped the stages of fact-checking process and the most urgent problems of fact-checkers (resulting from our validation survey) with the corresponding implications for AI support (Fig. 13).
Figure 13. Mapping of the top 10 most important problems of European fact-checkers with the corresponding stages of the fact-checking process, selected existing tools, and implications for AI support
@@
Our proposed implications for AI support come from the current state-of-the-art research and technical possibilities in the area of AI – machine learning (ML) and NLP (including the latest development of generative large language models), and also from available datasets and tools and are consistent with the work by Nakov et al. (2021a). Compared to Nakov et al. (2021a), we add three more AI tasks: 1) check-worthy document detection, 2) mapping of the existing fact-checks to additional (newly appearing) online content, and 3) fact-check summarization and personalization.
When discussing these implications, we aim to emphasize the need for human-centered AI systems (and a role that the current generative AI can play there) as well as reflect on the specific challenges and needs of the European fact-checkers identified in the previous sections. At the same time, we recognize that some of the most important problems do not require AI-based solutions. In such cases, the implementation of suitable tools represents mainly an engineering challenge (e.g., how to implement a tool that fact-checkers can use to insert ClaimReview schema into their fact-checks), which is out of the scope of this work.
@@ 6.1. Implications for Check-worthy document detection
The first AI task aims to support filtering the check-worthy documents (that is, news articles, blog posts, or even posts on social media) from monitored sources, which addresses the European fact-checkers’ problem of filtering relevant content for fact-checking from the monitored outlets. It can be defined as follows: given an input set of documents, detect such documents that are worth fact-checking, and thus they should receive fact-checkers’ attention (e.g., they contain factual claims, they are potentially impactful and harmful to society, etc.). This task can be addressed either as a classification (a document is or is not check-worthy) or a ranking problem (prioritizing the most check-worthy documents). It should be addressed with information available at document creation time (i.e., without relying on user feedback that appears later). Given that false information spreads faster than true information (Vosoughi et al., 2018), the impact of a fact-check can be limited once the post has already become viral. On the other hand, early fact-checks can limit the spread of false information and can thus serve as a way of pre-bunking, which has already been shown to be effective in existing studies (Roozenbeek et al., 2022).
@@
We argue that this task should generally precede the task of check-worthy claims detection (extraction) to limit and prioritize the amount of potential misinformation content the fact-checkers need to examine. Current systems (e.g., a fact-checking tool used within Meta’s Third-Party Fact-Checking Program) often prioritize the virality of a post as noted by the fact-checkers in our as well as in the previous studies (noa, 2020; Dierickx, 2022; Micallef et al., 2022). In research work, check-worthiness detection is originally related to claims (i.e., identification of particular sentences, typically in political debates). Nevertheless, starting in 2020, the CheckThat! Lab introduced the detection of check-worthy tweets (Hasanain and Elsayed, 2022). Although tweets are natively short texts containing just a few sentences, we would like to emphasize the distinction with check-worthy claim detection (see below), and we consider this as the first step towards check-worthy document detection. At the moment, we are not aware of any works on check-worthiness detection for longer pieces of text (e.g., for social media posts from other platforms that do not pose such strict length limitations, or even for whole news articles/blogs) or for multimodal content. Such extension of current research works and the construction of a suitable dataset represents the possible next step.
@@
For an AI solution to assist the fact-checkers, it needs to target sources relevant for them, consider criteria that are typically used by fact-checkers when they decide what to fact-check, and provide a means of justification/explanation of the selection of the check-worthy documents. The last two points can be facilitated by a related, more generic task of credibility assessment. Credibility signals (e.g., ones proposed by W3C Credible Web Community Group [^24] https://credweb.org/) would help fact-checkers (as well as other media professionals) to pre-screen the document and decide whether to proceed with the manual in-depth investigation to determine the necessity to fact-check it. Such a tool would speed up fact-checkers’ comprehension of the online content. To detect such indicators, a wide variety of techniques may be used, from a simple lookup in whitelists/blacklists, through the automatic check of predefined credibility criteria (e.g., presence of an author’s name, the known editorial board of a newspaper, etc.) up to advanced ML/NLP models. These would, for example, classify typographical and stylistic characteristics reflecting psychological features influencing the reader’s sentiment, detect the use of logical fallacies, or classify the leaning (bias) of the text (Da San Martino et al., 2020; Dhoju et al., 2019). With the recent advancements of generative AI, namely large language models (LLMs), and their uptake for evaluation of various aspects of texts, such as text quality (Hada et al., 2024) or level of personalization (Wang et al., 2023; Zugecova et al., 2025), the use of LLMs for credibility signals detection is also being researched, as discussed in a recent survey by Srba et al. (2024). One such example is the work of Leite et al. (2024) which employed LLMs to detect 19 signals, such as the presence of bias, use of polarizing language, etc. The detected check-worthiness of documents can serve as one of the indicators if the detection also considers their potential harmfulness or their expected impact on society (see, e.g., (Alam et al., 2021)), which are the aspects usually considered by the fact-checkers (see Fig. 5).
@@ 6.2. Implications for Check-worthy claims detection
While the task of check-worthy document detection (and the related task of credibility assessment/credibility signals detection) has attracted researchers’ attention only recently, the check-worthy claims detection has been addressed by the researchers for a longer period, and several established approaches, as well as datasets, already exist (see, e.g., (Hassan et al., 2017; Srba et al., 2024)). This task can be defined as follows: given an input sentence (or a set of multiple sentences, for example, from a political debate), detect such sentences that are worthy of fact-checking [^25] While the definition of a check-worthy claim differs across literature, it is usually a factual statement, that the general public would be interested in knowing, whether it is true or not.. Most of the existing approaches have so far focused on text using transformer-based language models, such as BERT (Devlin et al., 2019). However, in 2024 edition of the CLEF-2024 CheckThat! Lab task on check-worthy claims detection, the best-performing solutions were already using LLMs, including variations of LLaMA, Mistral, Mixtral, or GPT (Barrón-Cedeño et al., 2024), which were either fine-tuned or used just in an in-context learning setting (Hasanain et al., 2024). There is also a growing need for multimodal approaches, since claims can be accompanied by images, they can themselves appear in images (e.g., a screenshot of a social media post), or images can modify/change their meanings. There are already the first multimodal datasets for claim detection (see, e.g., (Zlatkova et al., 2019; Cheema et al., 2022)).
Similar to the detection of check-worthy documents, for check-worthy claim detection solutions to be human-centered (and useful to the fact-checkers), they need to be able to justify why a given claim is considered check-worthy. This is where the current generative AI approaches could help by providing understandable justifications beyond simple highlighting of tokens/words considered important by the detector.